[Hallc_running] [New Logentry] RC update for Hall C, Monday swing to Tuesday swing

mack at jlab.org mack at jlab.org
Tue Mar 10 19:50:02 EDT 2020

Logentry Text:
     Monday OWL changed to longitudinal target production at 05:46, nominal HMS/SHMS -3.5/-3.4 GeV data with 3He target. 

    Cascade failures following a valve closure kept beam off for most of Tuesday DAY shift. 

    After the Tuesday RC meeting, the 1 hour "Be window background run" was completed. This changes the dominant scatterers in the SHMS from "Be windows + 3He target" to "Be windows + N2 at atmospheric pressure". Ironically, Zhiwen says the rate increased, perhaps a teachable moment in how thin the 3He target is. 

    Sylvester thought the scraping had returned on the 3He cell. https://logbooks.jlab.org/entry/3801476 Xiaochao had us move the beam spot by another 0.5 mm in the negative X direction. (Done in 0.1 mm steps.) Need more statistics, but there seems to be no scraping now. https://logbooks.jlab.org/entry/3801594 Jingyi has updated the nominal beam position on the paper sheet on the wall. https://logbooks.jlab.org/entry/3801590 

    Resumed longitudinal target production at 05:46, nominal HMS/SHMS -3.5/-3.4 GeV data with 3He target. ONGOING AS OF TUE SWING 

    Let me know about any non-pivot work you'd like to take place during the March 13-20 down in Hall C. So far on the list we've got
        a 3He compressor checkout for the x > 1 experiment next year (Meekins/RadCon);
        a hopefully straightforward SHMS dipole PS relay replacement followed by testing (Lassiter/Segal);
        and an important SOLID detector test the scope of which is defined in the HClist http://devweb.acc.jlab.org/CSUEApps/hclist/task/1809 (Malace et al).

(The pivot work has been planned in great detail. I just want a list of anything 'else' so we don't trip over ourselves, and so the next RC can brace himself for a week of cat herding.)

    Don't forget the temporary instructions for shift members regarding SHMS dipole power supply : https://logbooks.jlab.org/entry/3800710 . 

    4 target polarization measurements since Monday SWING: 54.3%, 52.0%, 58.6%, and 57.6%, all L. The latter too-good-to-be-true numbers triggered an effort on day shift to adjust the heaters until good convection was re-established.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3801633
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200310/ce8f1b68/attachment.html>

More information about the Hallc_running mailing list